A cell using Zn/Zn+4 (aq, 1M) and H2 / H+1 (aq, 1M) generates a cell potential of + 0.67 V under standard conditions.


A. Write the half-cell reaction at the anode.

B. Write the half-cell reaction at the cathode.

C. Write the electrochemical cell notation.

D. Calculate the standard reduction potential at the anode.



1
Expert's answer
2021-04-22T03:03:36-0400

A. Zn - 2e = Zn2+

B. 2H+ + 2e = H2

C. Zn | Zn2+ || H+ | H2, Pt

D. Ecell = EH - EZn

EH (cathode) = 0V

EZn (anode) = EH - Ecell = 0 - 0.67V = -0.67V